OpenStudy (anonymous):

sine of x divided by one minus cosine of x + sine of x divided by one minus cosine of x = 2 csc x

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

Nnesha (nnesha):

\[\huge\rm \frac{ \sin(x) }{ 1-cosine(x) } +\frac{ \sin(x) }{ 1 -cosine(x) }\] what is the common denominator ?

OpenStudy (anonymous):

1-cosine(x)?

OpenStudy (anonymous):

but the second one is sinx/1+cosx

Nnesha (nnesha):

well the common denominator is 1-cosine(x) obviously both denominator are same u will get same numerator (sinx)\[\huge\rm \frac{ \sin(x) }{ 1-cosine(x) } \]

Nnesha (nnesha):

huh uh-oh are you sure check your question again

OpenStudy (anonymous):

yea i checked it 1+cosx

Nnesha (nnesha):

http://prntscr.com/7vfzly

Nnesha (nnesha):

\[\huge\rm \frac{ \sin(x) }{ 1-cosine(x) } +\frac{ \sin(x) }{ 1 +cosine(x) }\] so common denominator is (1-cosinex)(1+cosinex)

OpenStudy (anonymous):

but the picture of the equation in the book says 1+cosx

Nnesha (nnesha):

alright you typed it wrong

OpenStudy (anonymous):

it did that itself it copy pasted that way lol

Nnesha (nnesha):

really ? :o you didin't type it ?

Nnesha (nnesha):

shocking how that can translate :o :/

OpenStudy (anonymous):

lol no that was weird tho right?

Nnesha (nnesha):

\[\huge\rm \frac{ sinx(1+cosx) +sinx(1-cosx)}{ (1-cosx)(1+cosx)}\] common denominator is (1-cosx)(1+cosx) and multiply numerator of first fraction by denominator of 2nd fraction *and multiply numerator of 2nd fraction by denominator of first fraction

Nnesha (nnesha):

ye ....

OpenStudy (anonymous):

and then you cancel out?

Nnesha (nnesha):

no no NO! no then u will get the original question

Nnesha (nnesha):

distribute!

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

wait what

Nnesha (nnesha):

yes familiar withe the foil method ? apply that at the denominator

OpenStudy (anonymous):

oh okay so i leave the numerator alone righ tnow

OpenStudy (anonymous):

1-cos2x?

Nnesha (nnesha):

yep right now distribute both parentheses by sinx 't the numerator

OpenStudy (anonymous):

what? that last sentence you said threw me off i guess the wording did

Nnesha (nnesha):

\[\huge\rm \frac{ \color{reD}{sinx}(1+cosx) +\color{red}{sinx}(1-cosx)}{ (1-cosx)^2}\] distribute both parentheses by red sin x

OpenStudy (anonymous):

1sinx+sinxcosx+1sinx-sinxcosx

OpenStudy (anonymous):

2sinx/1-cos2x

Nnesha (nnesha):

and cos^2x equal to what ? do u remember sin/cos identities ?

OpenStudy (anonymous):

csc?

Nnesha (nnesha):

nope

Nnesha (nnesha):

special identity \[\huge\rm sin^2x + \cos^2 x = 1\] solve this for cos^2x

OpenStudy (anonymous):

some codes came out for when you said special identity nothing else?

Nnesha (nnesha):

sin^2x + cos^2x = 1 solve for cos^2 x

OpenStudy (anonymous):

cos^2x=1-sin^2x

Nnesha (nnesha):

yes right so replace cos^2x by 1-sin^2x

OpenStudy (anonymous):

okay

Nnesha (nnesha):

\[\huge\rm \frac{ \color{reD}{2sinx}}{ 1-(1-sin^2x)}\] distribute parentheses by negative one

OpenStudy (anonymous):

its coding

Nnesha (nnesha):

ahh 2sinx over 1-(1-sin^2x)

OpenStudy (anonymous):

sorry lol im being a pain

OpenStudy (anonymous):

2sinx/sin^2x

Nnesha (nnesha):

yes sin^2x is same as sinx times sin x so 2sin over sinx sin x simplfy

OpenStudy (anonymous):

1/sin2x

Nnesha (nnesha):

is it sin^2 x ?

OpenStudy (anonymous):

yea but with one over it

Nnesha (nnesha):

no what about 2 ? it's 2sinx over sinx times sin x

Nnesha (nnesha):